How much power does it take a power lifter to lift 102lbs 0.66m of the ground in 4.3 seconds?
Answer #1

Energy in lifting= mgh= 102*0.454*9.8*0.66=299.52 joules

Power= energy/time= 299.52/4.3= 69.65 watt.